OpenStudy (anonymous):

The height (in centimeters) of a candle is a linear function of the amount of time (in hours) it has been burning. When graphed, the function gives a line with a slope of -0.06 . Suppose that the height of the candle after 12 hours is 21.8 centimeters. What will be the height of the candle after 18 hours?

OpenStudy (mertsj):

y=-.06x+b Replace x with 12, y with 21.8 and find b Then you will have your equation and you can replace x with 18 to find y, the height of the candle after 18 hours.

OpenStudy (anonymous):

Let me try it. Thank you.

OpenStudy (anonymous):

first find an equation:\[y=mx+b\]where m is the slope and b is the y intercept and x and y are the x and y values we know that slope is -0.06 and we know y=21.8 at x=12 lets plug that in:\[21.8=-0.06(12)+b\]\[21.8=-0.72+b\]\[b=22.52\]now we can write an equation:\[y=-0.06x+22.52\]plug in the 18 for x to find the height :)

OpenStudy (anonymous):

thank you!!!!!!!!

OpenStudy (anonymous):

so 21.44 centimeters?

OpenStudy (anonymous):

yup :D